# Mixed-Strategy Nash Equilibrium for Crowd Navigation

### v2

@@ IV Mixed-strategy Nash equilibrium for social navigation
We now show that Algorithm 2 is guaranteed to converge to a global Nash equilibrium of a general-sum mixed-strategy game. Furthermore, the converged Nash equilibrium guarantees a lower-bounded reduction of the joint expected collision risk among all agents.
@@
(General-sum game) We define a $N$-player general-sum mixed-strategy game. The objective of player $i$ is:
@@
$p_{/i}(s)$ $=\overline{\sum_{j\in\mathcal{I}/i}}p_{j}(s)$ (13)
where $D(\cdot\|\cdot)$ is the Kullback-Leibler(KL)-divergence between two distributions. We assume the nominal mixed-strategies $p^{\prime}_{i}$ are given a priori and accessible by all players.
@@
Algorithm 2 always converges. When $p^{\prime}_{i}=p_{i}^{\prime},\forall i\in\mathcal{I}$, the converged mixed-strategies $(p_{1}^{*},\dots,p_{N}^{*})$ assemble a global Nash equilibrium of the general-sum game, meaning the following equation holds for any $i\in\mathcal{I}$:
$p_{i}^{*}$ $=\operatorname*{arg\,min}_{p_{i}}J_{i}(p_{1}^{*},\dots,p_{i},\dots,p_{N}^{*})$
@@
We leave details of the full proof in the appendix. The proof depends on the following theorem, which by itself reveals another important property of the Bayesian updating scheme. ∎
@@
The Bayesian posterior in two-agent social navigation (7) is the global minimum of the following optimization problem:
@@
$=p^{\prime}_{i}(s)\cdot\exp\Big{(}{-}\mathbb{E}_{p_{j}}[r](s)\Big{)}\cdot\exp(\lambda-1)$ (19)
@@
$=\eta\cdot p^{\prime}_{i}(s)\cdot z(s|p_{j})$ (23)
@@
Theorem 2 can be extended to the Bayesian posterior of multi-agent crowd navigation (11).
@@
The Bayesian posterior in multi-agent social navigation (11) global minimum of the following optimization problem:
@@
Follow the same proof of Theorem 2, with the distribution $p_{j}$ being replaced by $p_{/i}$. ∎∎
@@
The objective function (17) is the summation of two sub-objectives that represents the two most important behavioral properties in crowd navigation: collision avoidance and goal-orientated behavior. The first sub-objective $\mathbb{E}_{p,p_{i}}[r]$ evaluates the expected risk of collision between two agents. On the other hand, minimizing the second objective—the KL-divergence with respect to the nominal mixed-strategy—preserves the agent’s original intent during collision avoidance.
@@
Theorem 3 indicates that Algorithm 2 is conceptually similar to the commonly used iterative best response (IBR) scheme. However, IBR does not guarantee convergence, nor does it recover the global Nash equilibrium, while our iterative scheme is guaranteed to recover a global mixed-strategy Nash equilibrium.
On top of the guaranteed convergence to the mixed-strategy Nash equilibrium, we show that the converged mixed-strategies also guarantee a lower-bounded reduction in expected collision risk.
@@
The converged mixed-strategies $(p_{1}^{*},\dots,p_{N}^{*})$ from Algorithm 2 has a lower joint expected collision risk, compared to the nominal mixed-strategies. Furthermore, the reduction in expected collision risk is lower-bounded by the summation of KL-divergence between each agent’s converged mixed-strategy and the nominal mixed-strategy:
$\sum_{i=1}^{N}\sum_{j=i+1}^{N}\mathbb{E}_{p_{i}^{\prime},p_{j}^{\prime}}[r]-\sum_{i=1}^{N}\sum_{j=i+1}^{N}\mathbb{E}_{p_{i}^{*},p_{j}^{*}}[r]\geq\sum_{i=1}^{N}D(p_{i}^{*}\|p_{i}^{\prime})$ (26)
@@
Our model is highly generalizable for social navigation, since Algorithm 2, Theorem 1 and Theorem 4 do not depend on any assumption on the specific form of mixed-strategy and collision risk function, thus are compatible with arbitrary types of mixed-strategies and collision risk function. On the other hand, the game objective structure (12) is specifically for social navigation—minimizing the KL-divergence between the converged mixed-strategy from the nominal strategy does not necessarily make sense for other game-theoretic interactions.
Figure 4: Demonstration of learning temporal kernels from trajectory datasets and how they influence the Bayes’ rule Nash equilibrium. In each column of (a) and (b), we show from left to right: The snapshots from the original human [71] and mouse dataset [72]; Temporal correlation from the transformed variance dataset (45) and the optimal kernel after training; Trajectory samples from the nominal mixed-strategies characterized by the learned kernels; Trajectory samples from the Nash equilibrium mixed-strategies.
V Learning gaussian process model for nominal mixed-strategy
In Algorithm 2 we assume the nominal mixed-strategy of each agent is given. This section describes constructing mixed-strategy as a Gaussian process (GP). A Gaussian process is a Gaussian distribution over functions [73], in our case, trajectories. Similar to finite-dimensional multivariate Gaussian distributions, a GP can be characterized by a mean function $m(t)$ and a covariance kernel $K(t,t^{\prime})$:
$s(t)\sim\mathcal{GP}\big{(}m(t),K(t,t^{\prime})\big{)}$ (27)
Since we model trajectory as a mapping from time to agent state, the Gaussian processes in this work are temporal stochastic processes
Assumption 2.
Throughout the paper, we assume we have a symmetric and stationary kernel $K(t,t^{\prime})$:
[…]
We first discuss choosing the GP mean function for each agent’s nominal mixed-strategy. Here we construct the mean function for humans as a constant velocity model based on the current measured velocity [^2] As pointed out in [28], a constant velocity model could outperform state-of-the-art learning-based methods for short-term trajectory prediction. Especially since we assume nominal mixed-strategies do not encode collision behavior, the constant velocity model is a suitable model in our case.. For the robot, we use a high-level meta-planner to plan a path toward the goal as the mean of the nominal mixed-strategy. In an unstructured environment, the meta-planner can be the standard path planner (e.g., RRT) on top of the occupancy grid map of the environment.
With the GP mean functions constructed, the characterization of GPs as nominal mixed-strategies is equivalent to the characterization of GP kernel functions. Here, we introduce a method for learning GP kernel from offline trajectory datasets based on agents’ preference for maintaining a constant velocity. We extract this preference by measuring the temporal correlation across different times within a trajectory.
V-A Temporal correlation and conditional variance in GPs
We define temporal correlation as the correlation between the agent states across different times. Based on the marginal property of Gaussian processes, given a GP denoted as $\mathcal{GP}(m(t),K(t,t^{\prime}))$, we have that the marginal distribution of the GP at any two time steps $(t_{1},t_{2})$ is a finite-dimensional Gaussian distribution:
$\begin{bmatrix}s(t_{1})\\ s(t_{2})\end{bmatrix}$ $\sim\mathcal{N}\left(\begin{bmatrix}m(t_{1})\\ m(t_{2})\end{bmatrix},\begin{bmatrix}K(t_{1},t_{1})&K(t_{1},t_{2})\\ K(t_{2},t_{1})&K(t_{2},t_{2})\end{bmatrix}\right)$ (34)
$=\mathcal{N}\left(\begin{bmatrix}m_{1}\\ m_{2}\end{bmatrix},\begin{bmatrix}K_{11}&K_{12}\\ K_{21}&K_{22}\end{bmatrix}\right)$ (39)
The covariance matrix entries $K_{11},K_{22}$ determine the uncertainty of the agent state at the time $t_{1}$ and $t_{2}$, respectively. The off-diagonal entry $K_{12}=K_{21}$ determines the correlation between the agent states at time $t_{1}$ and at time $t_{2}$. Furthermore, based on the conditional property of multivariate Gaussian distribution, the Gaussian distribution of the state $x_{2}=s(t_{2})$ at time $t_{2}$, conditioned on the state $x_{1}=x(t_{1})$ at time $t_{1}$, can be represented through the off-diagonal entry $K_{12}=K_{21}$:
$x_{2}\sim\mathcal{N}\Big{(}m_{1}{+}K_{12}K_{22}^{-1}(x_{1}{-}m_{2}),K_{11}{-}K_{12}K_{22}^{-1}K_{21}\Big{)}$ (40)
We can generalize the above formula of variance of conditional Gaussian in finite-dimensional space to the infinite-dimensional space of Gaussian processes.
Definition 11.
(Conditional variance in GP) Given a temporal covariance kernel function $K(\cdot,\cdot)$, (40) is the conditional variance between two time steps $t_{1},t_{2}$:
$c(t_{1},t_{2})=K(t_{1},t_{1}){-}K(t_{1},t_{2})K(t_{2},t_{2})^{-1}K(t_{2},t_{1})$ (41)
Lemma 2.
Based on Assumption 2, the covariance kernel is stationary and symmetric, thus the conditional variance defined in (41) is stationary and symmetric:
[…]
Conditional variance (41) essentially transforms a high-dimensional probability distribution of trajectories (GP) to a low-dimensional space of displacement variance, enabling a way to extract a covariance kernel from offline datasets.
V-B Learning GP kernel from datasets
To characterize GP kernels from trajectory datasets, we can optimize the kernel parameter, such that the conditional variance of the GP as a function (41) optimally matches the conditional variance estimated from trajectory datasets.
Consider a trajectory dataset, denoted as $\mathcal{T}=\{s_{i}\}_{i\in\mathcal{I}}$, where agent $i$’s trajectory across $T$ time steps is a vector $s_{i}=[x_{i}^{[1]},\dots,x_{i}^{[T]}]\in\mathbb{R}^{T\times 2}$. Given a difference between two time steps $\delta t$, we can transform the original trajectory dataset $\mathcal{T}$ into a displacement dataset $\mathcal{D}$:
$\mathcal{D}_{\delta t}=\{\Delta x_{i}^{[t]}\},\quad\Delta x_{i}^{[t]}=|x_{i}^{[t+\delta t]}-x_{i}^{[t]}|$ (44)
The displacement dataset $\mathcal{D}_{\delta t}$ can transformed into a variance dataset $\mathcal{V}_{\delta t}$:
$\mathcal{V}_{\delta t}$ =$=\Big{\{}\nu_{i}^{\delta t}\Big{\}}_{i\in\mathcal{I}}$ (45)
$\nu_{i}^{\delta t}$ $=var\left(\Delta x_{i}^{[1]},\dots,\Delta x_{i}^{[T-\delta t]}\right)$ (46)
where $var(\dots)$ represents the variance of a set of scalar variables.
Given a kernel function $K_{\alpha}(t_{1},t_{2})$, parameterized by the parameter $\alpha$, we can optimize the kernel parameter with a set of the variance datasets $\mathcal{V}_{\delta t}$ across different time step differences $\delta t$, through the following least-squares optimization problem:
$\alpha^{*}$ $=\operatorname*{arg\,min}_{\alpha}\sum_{\delta t}\left(c_{\alpha}^{\delta t}-\overline{\sum}_{i\in\mathcal{I}}\nu_{i}^{\delta t}\right)^{2}$ (47)
$c_{\alpha}^{\delta t}$ $=K_{\alpha}(0,0){-}K_{\alpha}(0,\delta t)K_{\alpha}(0,0)^{-1}K_{\alpha}(0,\delta t)$ (48)
We anecdotally demonstrate the sufficiency of our GP learning method by characterizing GP kernels from a human trajectory dataset [71] and a mouse behavior dataset [72]. The results are shown in Figure 4, where our method sufficiently distinguishes the difference in crowd navigation patterns between humans and mice.
V-C Condition GP on prior knowledge
The use of GP as the nominal mixed-strategy has one extra advantage: it allows us to condition uncertainty at specific time steps. Assuming we are interested in evaluating the nominal mixed-strategy at a set of time points $\mathbf{t}=[t_{1},\dots,t_{n}]$, we can transform the infinite-dimensional GP to a finite-dimensional multi-variate Gaussian distribution:
$\mathbf{s}$ $=[s(t_{1})\dots s(t_{n})]$
$\sim\mathcal{N}\left(\underbrace{[m(t_{1})\dots m(t_{n}))]}_{\mathbf{m}},\underbrace{\begin{bmatrix}K(t_{1},t_{1})\dots K(t_{1},t_{n})\\ \vdots\quad\ddots\quad\vdots\\ K(t_{n},t_{1})\dots K(t_{n},t_{n})\end{bmatrix}}_{\mathbf{\Sigma}}\right)$ (52)
However, following the stationary kernel assumption, the uncertainty of each time point in the above finite-dimensional Gaussian distribution (52) are all identical:
$\mathbf{K}_{i,i}=\mathbf{K}_{j,j},\quad\forall i,j\in[1,\dots,N]$ (53)
This property is suboptimal in practice, because we may have prior knowledge about the agent’s state uncertainty at certain time points. For example, the state uncertainty at the current time point should be smaller than the state uncertainty $5$ seconds later, or the state uncertainty when being close to a potential goal location should be smaller than the state uncertainty far from the goal location. Gaussian processes provide a natural solution to incorporate such information into uncertainty quantification. Given a set of time points $\bar{\mathbf{t}}=[\bar{t}_{1},\dots,\bar{t}_{m}]$ and the variance of the state uncertainty at these time points $\sigma=[\sigma_{1},\dots,\sigma_{n}]$, the new finite dimensional Gaussian distribution after incorporating the prior uncertainty knowledge is:
$\mathbf{s}$ $\sim\mathcal{N}\left(\mathbf{m},\mathbf{\Sigma}^{*}\right)$ (54)
$\mathbf{\Sigma}^{*}$ $=\mathbf{\Sigma}-\mathbf{\Sigma}^{\prime}(\overline{\mathbf{\Sigma}}+diag(\sigma))^{-1}\mathbf{\Sigma}^{\prime\top}$ (55)
where
$\mathbf{\Sigma}^{\prime}$ $=\begin{bmatrix}K(t_{1},\bar{t}_{1})&\dots&K(t_{1},\bar{t}_{m})\\ \vdots&\ddots&\vdots\\ K(t_{n},\bar{t}_{1})&\dots&K(t_{n},\bar{t}_{m})\end{bmatrix}$ (59)
$\overline{\mathbf{\Sigma}}$ $=\begin{bmatrix}K(\bar{t}_{1},\bar{t}_{1})&\dots&K(\bar{t}_{1},\bar{t}_{n})\\ \vdots&\ddots&\vdots\\ K(\bar{t}_{n},\bar{t}_{1})&\dots&K(\bar{t}_{n},\bar{t}_{n})\end{bmatrix}$ (63)
$diag(\sigma)$ $=\begin{bmatrix}\sigma_{1}&&\\ &\ddots&\\ &&\sigma_{n}\end{bmatrix}$ (67)
In Figure 3(a,b) and Figure 4, we condition the GP prior at the beginning and the end of the planning horizon with small uncertainty.
Remark 7.
The formula for conditioning GP covariance (55) is identical to the GP posterior in GP regression [73]. However, we are not conducting GP regression here since we assume the GP mean is given as a constant velocity model. Here we use the covariance formula in GP regression to characterize the uncertainty of the nominal mixed-strategies.
VI Sampling from mixed-strategy Nash equilibrium
Algorithm 2 describes how to find a Nash equilibrium for mixed-strategy crowd navigation, and Section V describes how to construct nominal mixed-strategy for each agent. However, real-time computation of Algorithm 2 is challenging, given the hardness of evaluating the high-dimensional integrals. Here we introduce sampling-based approximations of mixed-strategy Nash equilibrium, for the purpose of real-time path planning. The key insight for the real-time sampling approximation is to leverage the connection between Algorithm 2 and Bayesian updating and Monte-Carlo integration.
VI-A Sample-based mixed-strategy representation
Instead of maintaining mixed-strategies as continuous probability density functions, we now represent the mixed-strategy as a set of samples. We denote the sample set for each agent’s mixed-strategy as:
$\mathbf{p}_{i}=\{\mathbf{s}_{i,j}\}_{j=1,\dots,M}$ (68)
where $i$ is the agent index and $j$ is the sample index (we assume each agent has $M$ samples).
With the samples, we can asymptotically approximate the conditional likelihood (8) through Monte-Carlo integration:
$z(s|p_{/i})$ $=\exp\Big{(}{-}\mathbb{E}_{p_{/i}}[r](s)\Big{)}$ (69)
$\approx z(s|\mathbf{p}_{/i})$ (70)
$=\exp\left[{-}\overline{\sum_{i\in\mathcal{I}/i}}\left(\frac{1}{M}\sum_{j=1}^{M}r(s,\mathbf{s}_{i,j})\right)\right]$ (71)
Based on Bayes’ rule, for any sample $\mathcal{s}$,
VI-B Sample-based Nash equilibrium approximation
The approximated conditional likelihood (71) serves as the ratio between the posterior and prior distribution:
$\frac{p_{i}(s)}{p^{\prime}_{i}(s)}\approx\eta\cdot\exp\left[{-}\overline{\sum_{i\in\mathcal{I}/i}}\left(\frac{1}{M}\sum_{j=1}^{M}r(s,\mathbf{s}_{i,j})\right)\right]$ (72)
Furthermore, the closed-form asymptotic approximation (71) can be also considered as a ratio between target distribution (posterior) and proposal distribution (prior). Assuming samples from the prior distribution are accessible, we can apply tools from statistical simulation to approximate the propagation of mixed-strategies in Algorithm 2 with sample-based representation. Below we introduce two sampling schemes, which approximate Algorithm 2 through rejection sampling [74] and importance sampling [75], respectively.
Rejection sampling
The outline of the rejection sampling scheme is as follows:
  1. [nolistsep]
  2. 1.
    Draw sample $\mathbf{s}$ from the proposal distribution $p^{\prime}$;
  3. 2.
    Compute the ratio of the likelihood of measuring the sample in target distribution $p$ and proposal distribution $w=\frac{p(\mathbf{s})}{p^{\prime}(\mathbf{s})}$, as described in (72);
  4. 3.
    Uniformly draw a scalar sample $u$ between 00 and $1$;
  5. 4.
    If $\gamma u<w$, accept the sample $\mathbf{s}$, otherwise repeat the process. Here the user-specified parameter $\gamma>1$ serves as a constant, finite bound on the likelihood ratio $\frac{p(\mathbf{s})}{p^{\prime}(\mathbf{s})}$;
  6. 5.
    If trajectory prediction is required, compute the mean of the converged samples as the trajectory prediction.
Following this outline, the application of rejection sampling to Algorithm 2 is straightforward, since the likelihood ratio between the proposal distribution (nominal mixed-strategy) and target distribution (Bayesian posterior mixed-strategy) is already given in (71). A detailed pseudocode for the rejection sampling-based BRNE algorithm can be found in the appendix.
Importance sampling
Different from rejection sampling, importance sampling does not generate new samples. Instead, it assigns weights to samples from the proposal distribution (nominal mixed-strategy) to represent the target distribution (Bayesian posterior mixed-strategy). Denoting the weight for sample $\mathbf{s}_{i,j}$ as $\mathbf{w}_{i,j}$, the approximated conditional likelihood formula (71), with weighted samples, can be reformulated as:
$z(s|\mathbf{p}_{/i})$ $=\exp\left[{-}\overline{\sum_{i\in\mathcal{I}/i}}\left(\frac{1}{M}\sum_{j=1}^{M}\mathbf{w}_{i,j}\cdot r(s,\mathbf{s}_{i,j})\right)\right]$ (73)
We now give the outline of the importance sampling variant below:
  1. [nolistsep]
  2. 1.
    Draw sample $\mathbf{s}$ from the proposal distribution $p^{\prime}$. Initialize the weight for each sample as $1$;
  3. 2.
    Update sample weight with (73);
  4. 3.
    Once all sample weights are updated, normalize weights such that the averaged weight remains $1$;
  5. 4.
    If trajectory prediction is required, compute the weighted mean of the converged samples as the trajectory prediction.
A detailed pseudocode for the importance sampling-based BRNE algorithm can be found in the appendix.
Comparison of sampling schemes
Compared to important sampling, rejection sampling better approximates the posterior belief, since it re-generates new samples during iterations. However, the computation of rejection sampling is less efficient compared to importance sampling, due to the cost of re-generating samples. In practice, we use rejection sampling for explicit examination of the posterior beliefs (e.g., Figure 3(a,b) are generated in this way) and importance sampling for real-time planning of the robot (e.g., all the navigation experiments in this paper are conducted this way).
VI-C Sampling-based crowd navigation
We now describe the overall pipeline for robot crowd navigation supported by mixed-strategy Nash equilibrium:
  1. [nolistsep]
  2. 1.
    The robot perception stack returns the robot’s current position and surrounding pedestrians’ current positions and velocities;
  3. 2.
    Given a planning horizon, the robot computes the constant velocity model and combines it with a pre-trained GP kernel to generate nominal mixed-strategies for the surrounding humans. At the same time, the robot uses a meta-planner to plan a collision-free path within the environment (without humans) and combines it with the pre-trained GP kernel as the nominal mixed-strategy for the robot;
  4. 3.
    The robot runs sampling-based approximation of Algorithm 2, which returns a set of samples from the mixed-strategy Nash equilibrium among the robot and the surrounding humans;
  5. 4.
    The robot computes optimal control signals that follows the mean of the Nash equilibrium mixed-strategy, which is represented as a set of sample trajectories;
  6. 5.
    The robot executes the optimal control signals until the next replanning cycle, then repeat the above steps.
The optimal control signals from the Nash equilibrium mixed-strategy can be done through trajectory optimization or optimal control methods. We name our crowd navigation method Bayes’ rule Nash equilibrium (BRNE) navigation.
One of the most significant differences between our method and the existing social navigation methods is our method serves as a “correction” scheme on top of existing navigation methods, by taking the planned trajectories from these navigation methods as the nominal strategy for the robot. This design choice makes our method more flexible and easier to integrate into existing systems, such as the Robot Operating System (ROS) navigation stack [76, 77].

### v4

@@ IV Mixed strategy Nash equilibrium for crowd navigation
We now show that Algorithm 2 guarantees the recovery of the global Nash equilibrium of a mixed strategy game. Furthermore, the converged Nash equilibrium guarantees a lower-bounded reduction of the joint expected collision risk among all agents.
IV-A Preliminaries on game theory
@@
(Pure strategy game) Given the strategy space $\mathcal{S}$, in a pure strategy general-sum game, each agent optimizes an individual objective function that depends on other agents’ (pure) strategies:
$s_{i}^{*}=\operatorname*{arg\,min}_{s_{i}}f_{i}(s_{1},\dots,s_{i},\dots,s_{N}).$ (13)
Based on the above definition, it is clear that each agent cannot optimize their individual objective function in isolation from other agents. Thus, the conventional optimality criteria for single-agent optimal decision-making no longer apply. Instead, Nash equilibrium [10, 11] is proposed to describe an equilibrium state where each agent’s strategy is optimal with respect to all other agents’ current strategies.
Definition 14.
(Global pure strategy Nash equilibrium) A set of (pure) strategies from all agents, denoted as $(s_{1}^{*},\dots,s_{N}^{*})$, reach the global Nash equilibrium if and only if the following equality holds for all agents:
$s_{i}^{*}=\operatorname*{arg\,min}_{s_{i}}f_{i}(s_{1}^{*},\dots,s_{i},\dots,s_{N}^{*}),\quad\forall i\in\{1,\dots,N\}.$ (14)
Remark 4.
A Nash equilibrium is local, as opposed to global, when (14) only holds for a local region within the strategy space.
Pure strategy describes the deterministic decisions of an agent. When decisions are uncertain, the game formula can be extended to mixed strategies.
Definition 15.
(Mixed strategy game) Given the strategy space $\mathcal{S}$, in a mixed strategy general-sum game, each agent optimizes an individual objective function that depends on other agents’ mixed strategies:
$p_{i}^{*}(s)=\operatorname*{arg\,min}_{p_{i}}J_{i}(p_{1},\dots,p_{i},\dots,p_{N}),$ (15)
where each mixed strategy $p_{i}(s)$ is a probability distribution over the strategy space $\mathcal{S}$.
Remark 5.
In practice, the individual objective function of a mixed strategy game is often, but not necessarily, formulated as the expected value of the pure strategy objective with respect to the mixed strategies:
$J_{i}(p_{1},\dots,p_{N})=\mathbb{E}_{p_{1},\dots,p_{N}}[f_{i}(s_{1},\dots,s_{N})].$ (16)
Similarly, the definition of Nash equilibrium can also be extended to mixed strategies:
Definition 16.
(Global mixed strategy Nash equilibrium) A set of mixed strategies from all agents, denoted as $(p_{1}^{*}(s),\dots,p_{N}^{*}(s))$, reach the global Nash equilibrium if and only if the following equality holds for all agents:
$p_{i}^{*}(s)=\operatorname*{arg\,min}_{p_{i}}J_{i}(p_{1}^{*},\dots,p_{i},\dots,p_{N}^{*}),\quad\forall i\in\{1,\dots,N\}.$ (17)
Based on the above definitions, we define the following $N$-player mixed strategy game for crowd navigation. The objective of player $i$ is:
@@
$\mathbb{E}_{p_{i},p_{/i}}[r]$ $=\sum_{j\in\mathcal{I}/i}\mathbb{E}_{p_{i},p_{j}}[r]$ (19)
where $D(\cdot\|\cdot)$ is the Kullback-Leibler(KL)-divergence between two distributions, $\mathcal{I}$ is the set of all agent indices, and $\mathcal{I}/i$ is the set of indices excluding index $i$. We assume the nominal mixed strategies $p^{\prime}_{i}$ are given a priori and accessible by all players.
We will now show that Algorithm 2 is guaranteed to converge to the global mixed strategy Nash equilibrium of this game.
IV-B Mixed strategy Nash equilibrium for crowd navigation
@@
The sequence of mixed strategies $\{(p_{1}^{[k]},\dots,p_{N}^{[k]})\}_{k}$ in Algorithm 2 converges to a limit point $(p_{1}^{*},\dots,p_{N}^{*})$ such that:
$\forall\epsilon>0,\exists K\in\mathbb{N},$
[…]
$F(p_{1},\dots,p_{N})=\sum_{i=1}^{N}\sum_{j=i+1}^{N}\mathbb{E}_{p_{i},p_{j}}[r]+\sum_{i=1}^{N}D(p_{i}\|p^{\prime}_{i}).$ (21)
The limit point is the global Nash equilibrium (17) of the mixed strategy game (18).
@@
We leave details of the full proof in the appendix. The proof depends on the following theorem, which reveals another important property of the Bayesian update scheme. ∎
@@
The Bayesian posterior in the two-agent iterative Bayesian update process (8) is the global minimum of the following optimization problem:
@@
$=p^{\prime}_{i}(s)\cdot\exp\Big{(}{-}\mathbb{E}_{p_{j}}[r](s)\Big{)}\cdot\exp(\lambda-1).$ (26)
@@
$=\eta\cdot p^{\prime}_{i}(s)\cdot z(s|p_{j}).$ (29)
@@
Remark 6.
The exponential-based weight update scheme, as well as its mathematical optimality, are not unique to crowd navigation and are not restricted to the formula in this paper. For example, in [79] the same weight update scheme is discussed in the context of online learning.
Theorem 2 can be extended to the multi-agent Bayesian update formula (12).
@@
The Bayesian posterior of the multi-agent Bayesian update formula (12) is the global minimum of the following optimization problem:
@@
Follow the same proof of Theorem 2, with the function $\mathbb{E}_{p_{j}}[r](s)$ being replaced by $\mathbb{E}_{p_{/i}}[r](s)$. ∎∎
@@
The individual objective function in our game formula (18) is the summation of two sub-objectives representing the two most important behavioral properties in crowd navigation: collision avoidance and goal-orientated behavior. The first sub-objective $\mathbb{E}_{p,p_{/i}}[r]$ evaluates the expected risk of collision between agents. On the other hand, minimizing the second objective—the KL-divergence with respect to the nominal mixed strategy—preserves the agent’s nominal intent during collision avoidance.
@@
Theorem 3 indicates that each Bayesian posterior (12) is the agent’s optimal response to the rest of the agents’ current mixed strategies. This means Algorithm 2 is equivalent to the commonly used iterative best response (IBR) scheme in game theory. However, IBR does not guarantee convergence, nor does it recover the global Nash equilibrium, while our iterative scheme is guaranteed to recover a global mixed strategy Nash equilibrium.
In addition to the guaranteed convergence, we show that the converged mixed strategies also guarantee a lower-bounded reduction in the expected collision risk.
@@
The converged mixed strategies $(p_{1}^{*},\dots,p_{N}^{*})$ from Algorithm 2 has a lower joint expected collision risk, compared to the nominal mixed strategies. Furthermore, the reduction in the expected collision risk is lower-bounded by the summation of KL-divergence between each agent’s converged mixed strategy and the nominal mixed strategy:
$\sum_{i=1}^{N}\sum_{j=i+1}^{N}\mathbb{E}_{p_{i}^{\prime},p_{j}^{\prime}}[r]-\sum_{i=1}^{N}\sum_{j=i+1}^{N}\mathbb{E}_{p_{i}^{*},p_{j}^{*}}[r]\geq\sum_{i=1}^{N}D(p_{i}^{*}\|p_{i}^{\prime}).$ (32)
@@
Our model is highly generalizable for crowd navigation. Algorithm 2, Theorem 1 and Theorem 4 do not depend on any assumption on the specific form of mixed strategy and collision risk function, thus are compatible with arbitrary types of mixed strategies and collision risk function. On the other hand, the game objective structure (18) is specifically for crowd navigation—minimizing the KL-divergence between the converged mixed strategy and the nominal strategy does not necessarily make sense for other game-theoretic interactions.
IV-C Connections to stochastic control and stochastic game-theoretic methods
Our framework consists of elements commonly found in other stochastic control and stochastic game-theoretic methods. Here, we discuss the connection between our framework and representative works from these fields.
Information-theoretic duality
The derivation of Theorem 2 and Theorem 3 is equivalent to the derivation of the information-theoretic duality between free energy and relative entropy (KL-divergence) in stochastic control [80]. In particular, if we specify the mixed strategy as a stochastic control process and specify the nominal mixed strategy as the path distribution of an uncontrolled system, then the individual objective function (18) is the upper bound of the agent’s free energy. In this case, Theorem 2 and Theorem 3 prove that each BRNE iteration globally minimizes the upper bound of each agent’s free energy, connecting our theoretical results to other results in [80].
Iterative best response and importance sampling
Our BRNE algorithm is based on two fundamental numerical techniques: iterative best response (IBR) and importance sampling. IBR is a widely used technique to efficiently approximate the Nash equilibrium of a game. Importance sampling is also widely used in stochastic control, particularly path integral control frameworks [73], to find the optimal control policy. In [61], a model predictive path integral control framework is combined with IBR (BR-MPPI) to find the optimal stochastic control policy for a vehicle interacting with other vehicles. While our BRNE framework and BR-MPPI share similar numerical techniques, there are two fundamental differences: (1) BR-MPPI is a general-purpose game-theoretic planner and BRNE is a game-theoretic crowd navigation planner; (2) Benefitting from its narrow scope, BRNE guarantees the convergence to the global Nash equilibrium and has sufficient computation efficiency to interact with a relatively large number of agents in crowds, while BR-MPPI does not have these advantages when applied to crowd navigation.
Bayesian update and multimodality
The stochastic formulation of BRNE is motivated by the uncertain nature of human behavior—one main advantage of the stochastic formula is that BRNE could capture the multimodality in an agent’s strategy (see Figure 2 for examples). Similarly, a stochastic game-theoretic planner (MPOGames) is proposed in [63], which explicitly computes multi-modal control policies in multi-agent interaction. The key difference between BRNE and MPOGames is the assumed origin of the multimodality in agent intents. The multimodality in MPOGames comes from the assumption that non-ego agents’ objectives are only partially observable to the ego agent. As a result, the ego agent maintains a belief over possible non-ego agent objectives using the Bayesian update, which creates multimodality. On the other hand, in BRNE, all agents’ objectives are assumed to be shared among all agents. The multimodality in BRNE comes from the global Nash equilibrium of a mixed strategy game recovered by BRNE, where the recovered Nash equilibrium could contain multimodal mixed strategies. Essentially, MPOGames and BRNE are making orthogonal contributions: MPOGames paves the way for extending BRNE to scenarios with incomplete information regarding agent objectives, and BRNE offers a potential solution to extend MPOGames beyond Gaussian-based policies and local approximation of Nash equilibrium.
V Evaluation